OpenStudy (anonymous):

If the parallel sides of a trapezoid contained by the lines y = -x + 5 and y = -x - 9, what equation represents the line contained by the midsegment?

OpenStudy (anonymous):

okay so what do you think

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

I'm not too sure, could you explain please? @Daniel56k

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know segment addition postulate

OpenStudy (anonymous):

sorry i don't understand what you just explained

OpenStudy (anonymous):



OpenStudy (anonymous):

dc=ab+ef

OpenStudy (anonymous):

meaning you have \[\frac{ (-x+5)+(-x-9) }{ 2 }\]

OpenStudy (anonymous):

can you try and finish it up

OpenStudy (anonymous):

I got -x - 2. Would that make the answer y = -x - 2?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you! @Daniel56k